ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 175, 2010

Aflalo, E. and J. Nekovář Non-triviality of CM points in ring class field towers

225

Aicardi, F. and L. G. Fel Gaps in nonsymmetric numerical semigroups

85

Amberg, B. and L. Kazarin On the product of two Chernikov subgroups

363

Aubert, A.-M., U. Onn, A. Prasad and A. Stasinski On cuspidal representations of general linear groups over discrete valuation rings

391

Bary-Soroker, L. and D. Kelmer On PAC extensions and scaled trace forms

113

Ben Simon, G. and D. A. Salamon Homogeneous quasimorphisms on the symplectic linear group

221

Borodzik, M. and H. Żoł _{}adek Complex algebraic plane curves via Poincaré–Hopf formula. II. Annuli

301

Choi, D. and Y. Choie p-adic limit of the Fourier coefficients of weakly holomorphic modular forms of half integral weight

61

Choie, Y. See D. Choi Chuang, C.-L., T.-K. Lee, C.-K. Liu and Y.-T. Tsai Higher

derivations of Ore extensions

157

Danilenko, A. I. and A. H. Dooley Simple ℤ^{2}-actions twisted by aperiodic automorphisms

285

Dolfi, S., E. Pacifici and L. Sanus Finite groups with real conjugacy classes of prime size

179

Dooley, A. H. See A. I. Danilenko

Evans, R. Seventh power moments of Kloosterman sums  349

Fathi, A. and A. Figalli Optimal transportation on non-compact manifolds

1

Fel, L. G. See F. Aicardi Figalli, A. See A. Fathi

Gitik, M. On normal precipitous ideals  191
Hatami, H. Graph norms and Sidorenko’s conjecture  125

Hinich, V. Plumbing coordinates on Teichmüller space: a counterexample

151

Kazarin, L. See B. Amberg Kelmer, D. See L. Bary-Soroker

Krupiński, K. Fields interpretable in rosy theories  421

Lee, T.-K. See C.-L. Chuang Liu, C.-K. See C.-L. Chuang Nekovář, J. See E. Aflalo Onn, U. See A.-M. Aubert Pacifici, E. See S. Dolfi Prasad, A. See A.-M. Aubert Salamon, D. A. See G. Ben Simon Sanus, L. See S. Dolfi Stasinski, A. See A.-M. Aubert Tsai, Y.-T. See C.-L. Chuang Żoł _{}adek, H. See M. Borodzik